OpenStudy (anonymous):

**Calc2 Help Needed** Find the area of the surface generated by revolving around the curve y=x^3 / 2 from x= 0 to 2 about the x-axis.

OpenStudy (anonymous):

\[2\pi \int\limits x^3/2 \sqrt ( (4+9x^4)/4\]

OpenStudy (dumbcow):

 to find surface area, we add up all the circumferences times dx \[SA = 2 \pi \int\limits_{a}^{b}R dx\]